# Differentially Private Linear Regression and Synthetic Data Generation with Statistical Guarantees

### v1

@@ 3.2 DP Binning-Aggregation Algorithms
$\tilde{t}_{k}$ $=t_{k}+\xi^{y},\quad\xi^{y}\sim\mathcal{N}(0,\Delta^{2}_{y}/\mu^{2}_{t}).$
@@ 4.2 Statistical Inference
As the number of bins $K\to\infty$ with the sample size $n\to\infty$, given the injected DP noise have finite variances, the bias-corrected estimator $\tilde{\bm{\beta}}\stackrel{{\scriptstyle def}}{{=}}(\widetilde{S}^{\top}\widetilde{W}\widetilde{S}-\widetilde{D})^{-1}\widetilde{S}^{\top}\widetilde{W}\widetilde{{\bm{t}}}$ as defined in Algorithm 2 satisfies
$\widetilde{\Sigma}^{-1/2}(\tilde{\bm{\beta}}-\bm{\beta})\xrightarrow{d}\mathcal{N}\left(0,\bm{I}_{d}\right),$
@@
$\widetilde{\Sigma}=\widetilde{M}^{-1}\widetilde{H}\widetilde{M}^{-1},$
with
$\widetilde{M}$ $=\frac{1}{K}(\widetilde{S}^{\top}\widetilde{W}\widetilde{S})-\widetilde{D}$,
$\widetilde{H}$ $=\frac{1}{K(K-d)}\sum_{k=1}^{K}\widetilde{\bm{Q}}_{k}\widetilde{\bm{Q}}_{k}^{\top},$
$\widetilde{\bm{Q}}_{k}$ $=\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\tilde{\bm{\beta}})+\tilde{w}_{k}D_{k}\tilde{\bm{\beta}}$.
Therefore, an asymptotic $(1-\alpha)$ confidence interval for each coefficient $\beta_{j}$ satisfying DP is given by
$\tilde{\bm{\beta}}_{j}\pm z_{\alpha/2}\cdot\mathrm{se}(\tilde{\bm{\beta}}_{j}),$
where $\mathrm{se}(\tilde{\bm{\beta}}_{j})$ is the $j$th diagonal entry of $\widetilde{\Sigma}$, and $z_{\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution.
@@ D Proofs
We prove the result by using the estimating equation theory.
Let $\xi_{{\bm{s}}_{k}}\sim\mathcal{N}(\bm{0},\sigma^{2}_{{\bm{s}}_{k}})$ and $\xi_{t_{k}}\sim\mathcal{N}(\bm{0},\sigma^{2}_{t})$ represent the Gaussian noise components that are added to the covariate sum vector ${\bm{s}}_{k}$ and the response sum variable $t_{k}$, respectively. Consequently, the observed noisy variables can be expressed as
@@
where the intrinsic error term is given by $\eta_{k}\sim\mathcal{N}(0,\sigma^{2}c_{k})$.
We define the bias-corrected estimator $\tilde{\bm{\beta}}$ as the solution to the estimating equation
$\bm{Q}(\bm{b})=\frac{1}{K}\sum_{k=1}^{K}\bm{Q}_{k}(\bm{b})=-\frac{1}{K}\sum_{k=1}^{K}\{\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\bm{b})+\tilde{w}_{k}D_{k}\bm{b}\}=\bm{0},$
where $D_{k}=\mathbb{E}(\xi_{{\bm{s}}_{k}}\xi_{{\bm{s}}_{k}}^{\top})=\text{diag}(\sigma^{2}_{\bm{s}_{k1}},\sigma^{2}_{\bm{s}_{k2}},\dots,\sigma^{2}_{\bm{s}_{kd}})$. At the true parameter value $\bm{\beta}$, the estimating equation is unbiased. For any index $k$, we have
$\mathbb{E}[\bm{Q}_{k}(\bm{\beta})]$ $=-\mathbb{E}\left[\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\beta)+\tilde{w}_{k}D_{k}\bm{\beta}\right]$
$=-\mathbb{E}\left[(\bm{s}_{k}+\xi_{{\bm{s}}_{k}})\tilde{w}_{k}(\bm{s}_{k}^{\top}\bm{\beta}+\eta_{k}+\xi_{t_{k}}-({\bm{s}}_{k}+\xi_{{\bm{s}}_{k}})^{\top}\bm{\beta})\right]-\mathbb{E}[\tilde{w}_{k}D_{k}\bm{\beta}]$
@@
$=\mathbb{E}[\tilde{w}_{k}]D_{k}\bm{\beta}-\mathbb{E}[\tilde{w}_{k}]D_{k}\bm{\beta}$
$=\bm{0}.$
That is, the estimating function is indeed unbiased at the true value $\bm{\beta}$. Furthermore, since the estimating equation is linear in the parameter vector $\bm{b}$, its Jacobian takes the following form:
$\widetilde{M}\stackrel{{\scriptstyle def}}{{=}}\frac{\partial\bm{Q}(\bm{b})}{\partial\bm{b}}=\frac{1}{K}\sum_{k=1}^{K}(\tilde{w}_{k}\tilde{\bm{s}}_{k}\tilde{\bm{s}}_{k}^{\top}-\tilde{w}_{k}D_{k})=\frac{1}{K}\widetilde{S}^{\top}\widetilde{W}\widetilde{S}-\widetilde{D}$,
where
$\widetilde{D}\stackrel{{\scriptstyle def}}{{=}}\frac{1}{K}\sum_{k=1}^{K}\tilde{w}_{k}D_{k}.$
Using a first-order Taylor expansion of the estimating function around the true parameter value $\bm{\beta}$, we obtain
@@
which can be rearranged to yield the following linear approximation:
@@
Let us now define
$H=\frac{1}{K^{2}}\sum_{k=1}^{K}\mathrm{Var}(\bm{Q}_{k}),$
where
$\bm{Q}_{k}\stackrel{{\scriptstyle def}}{{=}}\bm{Q}_{k}(\bm{\beta})=\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\bm{\beta})+\tilde{w}_{k}D_{k}\bm{\beta}.$
Then, by the classical Central Limit Theorem (CLT), as $K\to\infty$,
$H^{-1/2}\bm{Q}(\bm{\beta})\xrightarrow{d}\mathcal{N}\left(\bm{0},\bm{I}_{d}\right).$
A consistent estimator of $H$ can be obtained by its empirical counterpart, defined as
$\widetilde{H}=\frac{1}{K(K-d)}\sum_{k=1}^{K}\widetilde{\bm{Q}}_{k}\widetilde{\bm{Q}}_{k}^{\top},$
where
$\widetilde{\bm{Q}}_{k}=\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\tilde{\bm{\beta}})+\tilde{w}_{k}D_{k}\tilde{\bm{\beta}}$.
Finally, define the asymptotic covariance matrix estimator as
$\widetilde{\Sigma}=\widetilde{M}^{-1}\widetilde{H}\widetilde{M}^{-1}.$
Then, it follows that
$\widetilde{\Sigma}^{-1/2}(\tilde{\bm{\beta}}-\bm{\beta})\xrightarrow{d}\mathcal{N}(\bm{0},\bm{I}_{d}).$

### v2

@@ 3.2 DP Binning-Aggregation Algorithms
$\tilde{t}_{k}$ $=t_{k}+\xi^{y},\quad\xi^{y}\sim\mathcal{N}(0,B^{2}_{y}/\mu^{2}_{t}).$
@@ 4.2 Statistical Inference
As the number of bins $K\to\infty$ with the sample size $n\to\infty$, assume the injected DP noises have finite variances. The bias-corrected estimator
$\tilde{\bm{\beta}}:=(\widetilde{S}^{\top}\widetilde{W}\widetilde{S}-\widetilde{D})^{-1}\widetilde{S}^{\top}\widetilde{W}\widetilde{{\bm{t}}}$
satisfies
$\sqrt{K}\,(\tilde{\bm{\beta}}-\bm{\beta})\ \xrightarrow{d}\ \mathcal{N}\!\left(\bm{0},\;M^{-1}HM^{-1}\right),$
@@
$H:=\lim_{K\to\infty}\frac{1}{K}\sum_{k=1}^{K}\operatorname{Var}\!\big(\bm{Q}_{k}(\bm{\beta})\big),\quad\bm{Q}_{k}(\bm{\beta}):=\tilde{\bm{s}}_{k}\tilde{w}_{k}\big(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\bm{\beta}\big)+\tilde{w}_{k}D_{k}\,\bm{\beta},$
and
$\widetilde{M}:=\frac{1}{K}\Big(\widetilde{S}^{\top}\widetilde{W}\widetilde{S}\Big)-\widetilde{D}$,
with $\widetilde{M}\xrightarrow{p}M$ for some finite, nonsingular matrix $M$.
[…]
$\widetilde{\Sigma}=\frac{1}{K}\,\widetilde{M}^{-1}\widetilde{H}\,\widetilde{M}^{-1},$
where
$\widetilde{H}=\frac{1}{K-d}\sum_{k=1}^{K}\widetilde{\bm{Q}}_{k}\widetilde{\bm{Q}}_{k}^{\top},\qquad\widetilde{\bm{Q}}_{k}=\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\tilde{\bm{\beta}})+\tilde{w}_{k}D_{k}\,\tilde{\bm{\beta}}$.
Therefore, an asymptotic $(1-\alpha)$ confidence interval for $\beta_{j}$ is
$\tilde{\beta}_{j}\ \pm\ z_{\alpha/2}\,\sqrt{[\widetilde{\Sigma}]_{jj}},\qquad j=1,\dots,d.$
@@ D Proofs
The third result concerns the asymptotic distribution of the bias-corrected estimator $\tilde{\bm{\beta}}$. The argument follows from the general theory of estimating equations.
We begin by recalling the noise model. Let $\xi_{\bm{s}_{k}}\sim\mathcal{N}(\bm{0},\Sigma_{\bm{s}_{k}})$ denote the Gaussian noise added to the sufficient statistic $\bm{s}_{k}$, and let $\xi_{t_{k}}\sim\mathcal{N}(0,\sigma_{t}^{2})$ denote the noise added to $t_{k}$. Then the privatized quantities are
@@
where $\eta_{k}\sim\mathcal{N}(0,\sigma^{2}c_{k})$ is the regression noise.
The estimator $\tilde{\bm{\beta}}$ is defined implicitly as the solution to the estimating equation
$\bm{Q}(\bm{b})=\frac{1}{K}\sum_{k=1}^{K}\bm{Q}_{k}(\bm{b})=\bm{0},$
with
$\bm{Q}_{k}(\bm{b})=-\tilde{\bm{s}}_{k}\tilde{w}_{k}\bigl(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\bm{b}\bigr)-\tilde{w}_{k}D_{k}\bm{b},\qquad D_{k}=\mathbb{E}(\xi_{\bm{s}_{k}}\xi_{\bm{s}_{k}}^{\top}).$
At the true parameter $\bm{\beta}$, we check unbiasedness:
$\mathbb{E}[\bm{Q}_{k}(\bm{\beta})]$ $=-\mathbb{E}\Big[(\bm{s}_{k}+\xi_{\bm{s}_{k}})\tilde{w}_{k}\bigl(\bm{s}_{k}^{\top}\bm{\beta}+\eta_{k}+\xi_{t_{k}}-(\bm{s}_{k}+\xi_{\bm{s}_{k}})^{\top}\bm{\beta}\bigr)\Big]-\mathbb{E}[\tilde{w}_{k}D_{k}\bm{\beta}]$
@@
$=0,$
due to the independence between $\xi_{\bm{s}_{k}}$ and $\tilde{w}_{k}$. Hence, the estimating equation is unbiased.
We next analyze the asymptotic distribution. Because $\bm{Q}(\bm{b})$ is linear in $\bm{b}$, its Jacobian with respect to $\bm{b}$ is constant:
$\widetilde{M}=\frac{\partial\bm{Q}(\bm{b})}{\partial\bm{b}}=\frac{1}{K}\sum_{k=1}^{K}\bigl(\tilde{w}_{k}\tilde{\bm{s}}_{k}\tilde{\bm{s}}_{k}^{\top}-\tilde{w}_{k}D_{k}\bigr)=\frac{1}{K}\,\widetilde{S}^{\top}\widetilde{W}\widetilde{S}-\widetilde{D}$,
where $\widetilde{D}=\tfrac{1}{K}\sum_{k=1}^{K}\tilde{w}_{k}D_{k}$. A first-order Taylor expansion of $\bm{Q}(\cdot)$ around $\bm{\beta}$ yields
@@
which rearranges to
@@
Since the bins $\{B_{k}\}$ are disjoint, the regression errors $\{\eta_{k}\}$ are independent, and the DP noises $\{(\xi_{\bm{s}_{k}},\xi_{t_{k}})\}$ are generated independently across bins, the vectors $\{\bm{Q}_{k}(\bm{\beta})\}_{k=1}^{K}$ are independent with finite second moments, so the multivariate CLT applies. We then have
$\sqrt{K}\,\bm{Q}(\bm{\beta})=\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\bm{Q}_{k}(\bm{\beta})\ \xrightarrow{d}\ \mathcal{N}(\bm{0},H),\qquad H=\lim_{K\to\infty}\frac{1}{K}\sum_{k=1}^{K}\operatorname{Var}\!\big(\bm{Q}_{k}(\bm{\beta})\big).$
If $\widetilde{M}\xrightarrow{p}M$ with $M$ nonsingular, Slutsky’s theorem yields
$\sqrt{K}\,(\tilde{\bm{\beta}}-\bm{\beta})\ \xrightarrow{d}\ \mathcal{N}\!\big(\bm{0},\;M^{-1}HM^{-1}\big).$
A consistent estimator of $H$ is
$\widetilde{H}=\frac{1}{K-d}\sum_{k=1}^{K}\widetilde{\bm{Q}}_{k}\widetilde{\bm{Q}}_{k}^{\top},\quad\widetilde{\bm{Q}}_{k}=\tilde{\bm{s}}_{k}\tilde{w}_{k}(\tilde{t}_{k}-\tilde{\bm{s}}_{k}^{\top}\tilde{\bm{\beta}})+\tilde{w}_{k}D_{k}\tilde{\bm{\beta}}$,
[…]
$\widetilde{\Sigma}=\frac{1}{K}\,\widetilde{M}^{-1}\widetilde{H}\,\widetilde{M}^{-1}.$
This establishes asymptotic normality of the bias-corrected estimator with the stated covariance. ∎